The market demand for cigarettes is given by the following equation: Q = 2000 - 100P. The market supply of cigarettes is given by the following equation:Q = -100 + 200P . Quantity is defined as the number of packs, price is defined as price per pack.a. What is the market equilibrium price of cigarettes?

b. What is the market equilibrium quantity of cigarettes?

c. Graphically illustrate the supply curve, demand curve, and market equilibrium.

d. What is the price of cigarettes that consumers pay?

e. What is the price of cigarettes that producers receive?

f. In order to discourage smoking, now the government institutes a unit tax of $3 per pack that is levied on consumers.

i. What is the after-tax market equilibrium price of cigarettes?

ii. What is the after-tax market equilibrium quantity of cigarettes?

iii. Graphically illustrate the supply curve, demand curve, and market equilibrium.

iv. What is the price of cigarettes that consumers pay after the tax is imposed?

v. What is the price of cigarettes that producers receive after the tax is imposed?

vi. How much revenue does the tax generate for the government?

g. Assume that the revenue from the cigarette tax is spent on activities that benefit neither the consumers nor the producers of cigarettes. What is the economic incidence of the tax? What fraction of the tax is borne by consumers? What fraction is borne by producers? Briefly explain your reasoning.
Answer #1

(a)

In equilibrium, market demand equals market supply.

2000 - 100P = - 100 + 200P

300P = 2100

P = $7

(b)

When P = 7,

Q = 2000 - (100 x 7) = 2000 - 700

= 1300

(c)

From demand function, when Q = 0, P = 2000/100 = 20 (vertical intercept) & when P = 0, Q = 2000 (horizontal intercept).

From supply function, when Q = 0, P = 100/200 = 0.5 (vertical intercept).

In following graph, D0 and S0 are demand and supply curves intersecting at point A with equilibrium price P0 (= 7) and quantity Q0 (= 1300).

(d)

Price paid by consumers = Equilibrium price = $7

(e)

Price received by producers = Equilibrium price = $7

NOTE: As per Chegg Answering Policy, 1st 5 parts have been answered.